OpenStudy (anonymous):

does the equation x+4=x+4 have more than one real number solution?

OpenStudy (anonymous):

Yes, the equation does have more than one real number solution. In order to see this, consider to yourself a few examples of real numbers as possible solutions. For instance, x=1 is a real number. So, 1+4=1+4, which is true. So x=1 is a solution. But then, for x=2, we have 2+4=2+4, which is also true. So x=2 is also a solution. In fact, the equation x+4=x+4 has an infinite number of real numbers solution. Which means that no matter which real number you pick, the equation will hold true.